5. You deposit £10,000 in a London bank for a year when the interest rate on...

5.

You deposit £10,000 in a London bank for a year when the interest rate on pounds is

10 percent and the $/£ exchange rate is $1.50 per pound. A year later the $/£

exchange rate moves to $1.38 per pound. Calculate the dollar rate of return on your

deposit.
Answer #1

ANSWER:

Dollar appreciation = (exchange rate year later - exchange rate presently) / exchange rate presently = (1.38 - 1.50) / 1.50 = -0.12 / 1.50 = -0.08 or -8%

Interest paid by london bank = 10%

dollar rate of return on your deposit = interest paid by london bank + dollar appreciation = 10% - 8% = 2%

so the dollar rate of return on the deposit is 2%.